OpenStudy (anonymous):

1 5 6 - / 7 - 2 8 Find The Product Write As A Mixed Number

OpenStudy (anonymous):

\[\frac{ 14 }{ 5 }\times \frac{ 50 }{ 9 }\]

OpenStudy (anonymous):

\[15\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

i Got It But Im Going To Change The Question To A Division

OpenStudy (anonymous):

why you are doing this it's more easy

OpenStudy (jiteshmeghwal9):